OpenStudy (anonymous):

lim n-> infinity of n/(10^n)

OpenStudy (anonymous):

it's in indeterminate form so I used l'hospitals rule to get...

OpenStudy (anonymous):

1/ 10^n ln10

OpenStudy (anonymous):

where from here?

OpenStudy (anonymous):

why'd you take the ln ?

OpenStudy (anonymous):

it's zero,10^n grow up much more fast than n

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

o wait nvm i thought it as msething else

OpenStudy (anonymous):

the derivative of the denominator is 10^n ln10

OpenStudy (anonymous):

yes simply put infinity in and you'll see that the bottom exponentialy grows while the top stays the same

OpenStudy (anonymous):

one is a line and other exponential,so it is zero

OpenStudy (anonymous):

1/infinity is zero...thanks!

OpenStudy (anonymous):

You can just notice that the bottom increases at a much faster rate than the numerator, so it must approach 0

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

\[\frac{d}{dn}[a^n]=ln(a)a^nn'=ln(10)10^n(1)\]

OpenStudy (anonymous):

sorry i'm just looking at your work weird because of it being on one line lol

OpenStudy (anonymous):

i see..lol